A fertilizer mixture containing 30% nitrogen is desired for a certain crop. The farmer has a fertilizer with 17% nitrogen content and another fertilizer with 43% nitrogen content. How much of each fertilizer is needed to make 100kg of fertilizer with the desired nitrogen content ?
please give me the table of your solution

Answer by checkley77(12844)

.43x+.17(100-x)=.3*100
.43x+17-.17x=30
.26x=30-17
.26x=13
x=13/.26
x=50 kg of 43% fertilizer is used.
100=50=50 kg of 17% fertilizer is also used.
Proof:
.43*50+.17*50=30
21.5+8.5=30
30=30